

********************************************************************
UARS WEEKLY STATUS REPORT - 06 AUG 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM gathered its normal solar, offset, and occultation data this past
week without apparent problems.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
HRDI is operating normally.
***********
WINDII
***********
WINDII is next scheduled for activation August 12-15, 1999.
Gordon Shepherd gave two WINDII presentations at the IUGG
meeting in Birmingham, U.K., July 19 - July 30.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
16 days CLA (7.0) PROD (L1-L3) (473R-158C-12F=303L)
11 days GPI (4.05) PROD (CLAES L3B) (69R-45C=24L)
9 days WIN (5.11) PROD (L1-L3) (1540R-1028C-33F=479L)
The CDHF completed the upgrade of Multinet to V4.2 as part of
the ongoing Y2K implementation effort.
Additional ISO compliant procedures for the new Alpha 4000's
(CDHFA/B - REHOST) are currently being drafted. These additions
reflect the new Storage Silo (HSM). Current CDHF User interfaces,
menus etc., will not change.
DAAC data created in the month of May 1999 has been loaded into the
DAAC transfer tables for HALOE, PEM, NMC, and UK
***********
FOT
***********
REPORTING PERIOD
This report covers 24 July 1999 (Orbit 42983, GMT Day 205) through 30 July 1999
(GMT Day 211).
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 206/0725 43002 Off
ACRIM 207/0013 43012 On
SUSIM 207/0013 43012 On
WINDII 207/0909 43018 On
MLS 208/2342 43042 Off
WINDII 209/2257 43056 Off
The beta angle ranged between the angles of 50.4 degrees to 74.1 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
75 degrees on 31 July 1999 (DOY 212).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.6 to 0.8 degrees C. Additional battery performance data for each
battery on 23 July 1999 and 30 July 1999 are provided in Appendix D.
The clock error ranged from -11.0 to 11.2 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 03
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 04
% WINDII = 02
% MLS = 02
% HRDI = 02
% HALOE = 01
% ACRIM = 01
% SUSIM = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 05
Four (4) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
% Transponder A: 99/179: 2287.496941
% Transponder B: 99/180: 2287.498977
GROUND SYSTEM OPERATION
Five (5) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 99/210 show the observatory data loss to be 24 hours, 54
minutes, 39.732 seconds (an increase of 1.088 seconds since the last report
period). This is a 0.0362 percent data loss which equals a 99.9638 percent
data capture for the mission.
The increase of 1.088 seconds occurred over one (1) day:
7 1.088 seconds on Day 206 due to no overlap.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT personnel are supporting the IMOC (SCS-21) development effort. The current
work is a continuation of the SCS-21 UARS implementation effort done last year
which culminated in a level 1 demonstration.
The FOT continues to investigate the feasibility of a new PMON action to change
battery V/T curves based on loading to mitigate battery risks in sun pointing
safehold.
YEAR 2000 OPERABILITY SUPPORT
The FOT received a set of Y2K planning aids from FDF which will be used to
facilitate UARS Y2K end-to-end testing. Eleven 3-day planning aid files were
received and processed by the FOT. Two of these files did not process
successful on the CMS due to either incorrect data, or invalid data formats.
FDF has been notified of the problems, and are in the process of correcting the
files.
MEETINGS
None.
FUTURE SIGNIFICANT EVENTS
August 20, 1999 (DOY 232): The next UARS Yaw Around Maneuver (forward-to
reverse) is scheduled.
Aug 11, 1999 (DOY 223): Preparations for an eclipse of the Sun is planned. All
instruments will be scheduled off. Final details will be available next week.
PERSONNEL CHANGES
None.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 30 July 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
42995 205/2008 99-252 MLS SCAN RESET
43002 206/0725 99-253 HALOE Power OFF
43011 206/2242 99-254 OBC Clk Adj 20.6 to 0.00 msec/day
43012 207/0013 99-255 ACRIM Power ON
43012 207/0013 99-256 SUSIM Power ON
43018 207/0909 99-257 WINDII Power ON
43027 207/2338 99-258 OBC Clk Adj 0.00 to 20.6 msec/day
43033 208/0909 99-259 OBC Clk Adj 20.6 to 41.2 msec/day
43039 208/1850 99-260 HRDI HRDI BARREL
43042 208/2342 99-261 MLS Power OFF
43052 209/1531 99-262 HRDI HRDI MPL
43056 209/2257 99-263 WINDII Power OFF
43067 210/1532 99-264 OBC Clk Adj fr 41.2 TO 20.6 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 30 July 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSBARTMP Y-LO HRDI BARREL TEMP CONTINUOUS
HRPTSBARTMP Y-HI HRDI BARREL TEMP 43043-45
HRPTSBARTMP R-HI HRDI BARREL TEMP 43046-56
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
No Alerts to report this report period.
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 30 July 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT
TIME
CLOCK RATE CHANGE
43011
206/22:52 (25 July 99)
from 20.6 to 00.0 msec/day
43027
207/23:44 (26 July 99)
from 00.0 to 20.6 msec/day
43033
208/09:15 (27 July 99)
from 20.6 to 41.2 msec/day
43067
210/15:35 (29 July 99)
from 41.2 to 20.6 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 30 July 1999
23 July 1999 - GMT Day 204 Beta = 50.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ----- ------- ------
1 -106.4/-134.4 -3.05/-4.62 +0.0/+0.0 16.5 0.000 0.0
2 +44.8/-22.4 +5.47/+4.19 +26.4/-16.0 25.1 1.021 21.0
3 +22.4/-50.4 +4.51/+3.23 +25.6/-15.6 25.0 1.023 20.7
30 July 1999 - GMT Day 211 Beta = 74.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ----- ------- ------
1 -112.0/-134.4 -2.42/-3.68 +0.0/+0.0 16.5 0.000 0.0
2 +22.4/-5.6 +2.60/+1.66 +16.4/-13.6 27.5 1.036 13.4
3 +22.4/-22.4 +1.97/+1.03 +16.0/-13.6 27.5 1.035 13.1
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 30 July 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-035 FOT 43006 206/1331 TAC Crashed
99-033 FOT 43030 208/0456 14170 Generic Late Acquisition
99-033 FOT 43032 208/0731 14170 Generic Late Acquisition
99-036 FOT 43050 209/1222 TAC Crashed
99-037 FOT 43055 209/2020 TAC Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-033 FOT see list July 99 14170 Generic Late Acq.
99-035 FOT 43006 206/1331 TAC Crashed
99-036 FOT 43050 209/1222 TAC Crashed
99-037 FOT 43055 209/2020 TAC Crashed
SEVNTFW99/31